# Multifidelity linear regression for scientific machine learning from scarce data

### v1

@@ 4 Numerical results
In this section, we present numerical results that demonstrate the efficacy of the proposed multifidelity linear regression method applied to an analytical example in Section 4.1 and a PDE model problem in Section 4.2. We compare models learned with our proposed multifidelity linear regression approach to models trained in the standard way on only high-fidelity data. To ensure fair comparisons, we compare models with equivalent training cost.
@@ 4.1 Analytic example: exponential function
$f^{(1)}(z)$ $=8\exp{z},$ (25)
$f^{(2)}(z)$ $=0.9\sqrt{8f^{(1)}(z)}=0.9\times 8\exp(0.5z),$ (26)
where $z\sim\mathcal{U}(0,5)$ with assumed costs of $w_{1}=1$ and $w_{2}=0.001$. For this analytical example, the correlation coefficient between the models is $\rho_{12}=0.97$. We fit a quadratic model to approximate $f^{(1)}(z)$, which leads to learning three regression coefficients for the one-dimensional problem as $\hat{f}(z;\beta)=\beta_{1}+\beta_{2}z+\beta_{3}z^{2}=x(z)^{\top}\beta$ where $x(z)=[1,z,z^{2}]^{\top}$. We show results for both the standard high-fidelity only and the proposed multifidelity approaches for three different computational budgets of 10, 100, and 1000.
Experiments using exact model statistics. For this analytical example, the exact model statistics $\sigma_{i}$, $\rho_{12}$, and $C_{12}$ and $C_{22}$ can be calculated. For a given budget, we use the exact statistics to compute the sample allocations according to eq. 10. The resulting sample allocation is shown in Table 1.
@@
Using this allocation, we compute multifidelity linear regression estimators using the three different strategies for choosing the control variate coefficient $\alpha$ or $A$ from Section 3.2. We compare the performance of these three multifidelity estimators in the setting where exact model statistics are known in order to validate the theoretical results from Section 3.3. The scalar control variate coefficient chosen by the MF-mean strategy (9) is $\alpha=12.79$. The MF-$\alpha^{*}$ optimal scalar control variate coefficient is $\alpha^{*}=11.70$. The MF-$A^{*}$ optimal matrix control variate coefficient is
$A^{*}=\begin{pmatrix}19.6&-7.3&1.3\\ 168.6&-69.9&10.5\\ 1330.6&-553.6&75.2\end{pmatrix}.$
The regression fits are repeated over 500 repetitions of training data to analyze robustness of the proposed method. Figure 1 plots realizations of the fitted regression models $\hat{f}(z;\hat{\beta})$ from both the standard and multifidelity training approaches for a computational budget of 100, together with a comparison to the exact regression model $\hat{f}(z;\beta^{*})$, which can be computed for this analytical example. We note that the realizations of both standard and multifidelity regression models are centered around the exact regression model, but the standard approach exhibits higher variance and is less robust to variations of the training data, even with the relatively high budget of 100 high-fidelity samples for this simple example.
Figure 1: Comparison of standard linear regression and multifidelity linear regression fits over 500 repetitions of training data for the analytic example using a computational budget of 100.
Figure 2 plots 500 realizations each of the first element of $\hat{C}_{XY}$, the first regression coefficient $\hat{\beta}_{1}$, and the regression model prediction at $z=5$, $\hat{f}(z=5;\hat{\beta})$, for estimators obtained via the standard single-fidelity as well as via our proposed multifidelity regression approaches for the three computational budgets 10, 100, and 1000. The realizations are plotted with some transparency so that regions with more realizations appear more densely colored. The mean of the realizations is plotted in the solid colored line on each plot, and the exact value (computable for this analytical example) is plotted in the dotted black line. These numerical results support the statement in Theorem 3.1 that both the single- and multifidelity estimators are unbiased; i.e., the estimator mean is the true statistic, regardless of the choice of control variate coefficient.
Figure 2 also illustrates the central thesis of this work: that multifidelity training leads to lower variance of the linear regression estimators and thus a more robust model, particularly in the low-budget regime. This is true for all three choices of control variate coefficient. The variance of the standard vs. multifidelity regression estimators is quantified in Figure 2(a), which plots the trace of the covariance of 500 realizations of estimators for $\hat{C}_{XY}$, of $\hat{\beta}$, and of the regression model predictions at $z=5$. As predicted by the theoretical analysis of Section 3.3, in this setting where the exact model statistics are known, the optimal matrix coefficient leads to the greatest variance reduction (Theorem 3.8), whereas the optimal scalar coefficient (Theorem 3.4) leads to a greater reduction than the sub-optimal scalar coefficient chosen according to (9). This variance reduction means that multifidelity learned models trained with the lowest budget (just 8 high-fidelity samples) have similar average error to standard learned models trained with hundreds of high-fidelity samples, again demonstrating the efficacy of the multifidelity approach for learning robust and accurate models in the scarce data regime.
@@
Figure 2: Analytical example: 500 realizations (semi-transparent colored lines) of estimators for first entry of $\hat{C}_{XY}$ (top row), first regression coefficient (middle row), and regression model prediction $\hat{f}(z=5;\hat{\beta})$, when true model statistics are known. Black dashed line denotes known true values and thick solid colored lines show the mean over realizations computed from different training data.
Experiments using inexact model statistics. Figures 2(b) and 2(c) plot the trace of the covariance for the 500 realizations of the standard and multifidelity regression estimators when the model statistics are estimated inexactly using pilot samples for each realization. Figure 2(b) uses a moderate pilot sample size of 100 whereas Figure 2(c) uses a small pilot sample size of 10. We find that using estimated model statistics to determine the sample allocation according to (10) yields a 1% (5%) variation in $m_{1}$ and a 5% (20%) variation in $m_{2}$ when 100 (10) pilot samples are used. Despite this variation in sample allocation, the multifidelity regression realizations have lower variance than the standard single-fidelity approach.
For this analytical example, the MF-$A^{*}$ strategy yields the greatest variance reduction even when it is computed from inexact model statistics, although the difference between its performance and that of the scalar coefficient choices becomes smaller as the model statistic estimate becomes poorer. The two choices of of scalar coefficient, MF-$\alpha^{*}$ and MF-mean, perform similarly when inexact model statistics are used in their calculation: the theoretically optimal MF-$\alpha^{*}$ barely outperforms the sub-optimal MF-mean strategy when 100 pilot samples are used, and sometimes underperforms when just 10 pilot samples are available. The good performance of the simpler MF-mean strategy in practical settings where model statistics must be estimated leads us to recommend it over the theoretically more optimal options that require estimations of more complex statistics.
(a) True model statistics
(b) 100 pilot samples
(c) 10 pilot samples
Figure 3: Comparing the effect of optimal control variate coefficients with respect to the minimization of the (generalized) variance of the estimators as described in Theorems 3.4 and 3.8 when (a) true model statistics are used compared to when model statistics are estimated from (b) 100 pilot samples and (c) 10 pilot samples for the three options proposed in Section 3.2.1 for the analytic example.
@@ 4.2.1 Convection-diffusion-reaction (CDR) problem
$\begin{split}\frac{\partial u}{\partial t}&=\kappa\Delta u-U\nabla u+s(u,z)\\ s_{i}(u,z)&=\nu_{i}\left(\frac{W_{i}}{\rho_{\text{mix}}}\right)\left(\frac{\rho_{\text{mix}}Y_{F}}{W_{F}}\right)^{\nu_{F}}\left(\frac{\rho_{\text{mix}}Y_{O}}{W_{O}}\right)^{\nu_{O}}A_{\text{pe}}\exp\left(-\frac{E}{RT}\right),\,\quad\text{for }i=F,O,P,\\ s_{T}(u,z)&=s_{P}(u,z)Q.\end{split}$ (27)
In (27), the thermo-chemical state is $u(z,t)=[Y_{F},Y_{O},Y_{P},T]^{\top}$, where $Y_{F}$, $Y_{O}$, and $Y_{P}$ are the mass fractions of the fuel, oxidizer, and product, respectively, and $T$ is the temperature; $\kappa$ is the molecular diffusivity, $U$ is the velocity field, and $s$ is the nonlinear reaction source term. The reaction modeled is a one-step hydrogen combustion given by $2\text{H}_{2}+\text{O}_{2}\rightarrow 2\text{H}_{2}\text{O}$, where hydrogen is the fuel, oxygen acts as the oxidizer, and water is the product. The source terms are defined using $A_{\text{pe}}$ as the pre-exponential factor of the Arrhenius equation, $E$ as the activation energy, $W_{i}$ as the molecular weight of species $i$, $\rho_{\text{mix}}$ as the density of the mixture, $R$ as the universal gas constant, and $Q$ as the heat of the reaction.
@@
where $T_{i}$ is the temperature at the inlet, $T_{0}$ is the temperature of the left wall of the domain, and $\phi$ is the fuel:oxidizer ratio of the premixed inflow. The parameters $A_{\text{pe}}$ and $E$ are assumed to be log-uniformly distributed and $T_{i}$, $T_{0}$, and $\phi$ are assumed to be uniformly distributed. We fit a quadratic model to these five inputs, so that the feature vector $x$ has length 21 (one constant, five linear, and 15 quadratic features).
@@ 4.2.2 Results: CDR problem
Results using high-sample model statistics estimates. Because analytical model statistics are not easily obtainable for this example, we first provide multifidelity regression results that use model statistics estimated from $10^{5}$ samples. For computational budgets of 10, 100, and 1000, these estimated model statistics yield the sample allocations tabulated in Table 3. Figure 4 then plots 500 realizations of the estimators for the first element of $\hat{C}_{XY}$, the first regression coefficient $\hat{\beta}_{1}$, and the learned model prediction at $z=a$ where $a$ is a given point in the input domain. The realizations are again plotted with some transparency; the mean of the estimator realizations is again shown in the solid colored line; and the estimators resulting from using $10^{5}$ high-fidelity samples is plotted using a dashed black line for reference. Figure 4 illustrates our theoretical results from Section 3: that the multifidelity estimators are unbiased (Theorem 3.1), and that when good estimates of the model statistics are known, the MF-$A^{*}$ optimal matrix coefficient $A^{*}$ yields greater variance reduction than the MF-$\alpha^{*}$ optimal scalar coefficient $\alpha^{*}$, which in turn yields greater variance reduction than the sub-optimal MF-mean $\alpha$.
@@
Table 3: Sample allocation based on (10) for the CDR problem when high-sample estimates of model statistics are available.
Figure 4: Illustrating unbiasedness of standard linear regression and multifidelity linear regression fits over 500 repetitions of training data for estimation of first entry of $\hat{C}_{XY}$ and regression coefficient along with the prediction $\hat{f}(z=a;\hat{\beta})$ where $a=[5.5\times 10^{11},6000,300,925,1]$ when high-sample estimates of model statistics are used for the CDR problem. Black dashed line denotes known true values and colored lines denote the estimates from 500 repetitions of training data.
Results using low-sample model statistic estimates. In Figure 5, we study the effect of using smaller pilot sample sizes to estimate the model statistics used to determine control variate coefficients and sample allocation. Figure 4(a) plots the (generalized) variance of $\hat{C}_{XY}$, $\hat{\beta}$, and $\hat{f}(z;\hat{\beta})$ when $10^{5}$ pilot samples are used, for reference. The generalized variance of the estimators is computed from 500 realizations of the standard/multifidelity linear regression estimators. Figure 4(b) shows the variance of the estimators when 100 pilot samples are used, and Figure 4(c) the results for just 10 pilot samples. We find that using estimated model statistics to determine the sample allocation according to (10) yields a 1% (10%) variation in $m_{1}$ and a 6% (25%) variation in $m_{2}$ when 100 (10) pilot samples are used. Despite this variation in sample allocation, the multifidelity regression realizations have lower variance than the standard single-fidelity approach.
For this PDE model problem, the gains from using MF-$A^{*}$ are not as significant as for the analytic example. In fact, for the smallest pilot sample size of 10 (Figure 4(c)), we see that MF-$A^{*}$ performs significantly worse than the scalar control variate coefficients. This is because the optimal $A^{*}$ choice depends on accurate estimation of the covariances $C_{12}$ and $C_{22}$ which are in $\mathbb{R}^{21\times 21}$ for this problem, and are thus difficult to estimate. This indicates that for complex engineering problems, the theoretically optimal MF-$A^{*}$ strategy should not be used when statistics must be estimated using limited pilot samples. However, multifidelity regressions using both choices of scalar coefficient perform well across different pilot sample sizes, yielding significant variance reduction when compared to standard linear regression for varying computational budgets. The optimal $\alpha^{*}$ performs better than the optimal $A^{*}$ because the choice of $\alpha^{*}$ relies only on accurate estimation of $\mathsf{Tr}(C_{12})$ and $\mathsf{Tr}(C_{22})$, which are scalar quantities which are easier to estimate from limited data. For this case, both scalar $\alpha$ multifidelity learned models trained with the lowest computational budget (just 4 high-fidelity samples) achieve a similar average error to standard learned models trained with $\mathcal{O}(100)$ high-fidelity samples, again demonstrating the efficacy of the multifidelity approach for learning robust and accurate models in the scarce data regime.
(a) High-sample model statistics using $10^{5}$ samples
(b) 100 pilot samples
(c) 10 pilot samples
Figure 5: Comparing the effect of optimal control variate coefficients with respect to the minimization of trace of the covariance of the estimators as described in Theorems 3.4 and 3.8 when model statistics are (a) high-sample estimates using $10^{5}$ samples compared to model statistics estimated from (b) 100 pilot samples and (c) 10 pilot samples for the three options proposed in Section 3.2.1 for the CDR problem.

### v2

@@ 4 Numerical results
In this section, we present numerical results that demonstrate the efficacy of the proposed multifidelity linear regression method applied to an analytical example in Section 4.1 and a PDE model problem in Section 4.2. We compare models learned with our proposed multifidelity linear regression approach (MF) to models trained in the standard way on only high-fidelity data (HF). To ensure fair comparisons, we compare models with equivalent training cost.
@@ 4.1 Analytic example: exponential function
$f^{(1)}(z)$ $=\exp{z},$ (20)
$f^{(2)}(z)$ $=0.9\sqrt{f^{(1)}(z)}=0.9\times\exp(0.5z),$ (21)
where $z\sim\mathcal{U}(0,5)$ with assumed costs of $w_{1}=1$ and $w_{2}=0.001$. For this analytical example, the correlation coefficient between the models is $\rho_{12}=0.97$. We fit a fourth-order polynomial model to approximate $f^{(1)}(z)$, which leads to learning five regression coefficients for the one-dimensional problem. That is, $\hat{f}(z;\beta)=\beta_{1}+\beta_{2}z+\beta_{3}z^{2}+\beta_{4}z^{3}+\beta_{5}z^{4}=x(z)^{\top}\beta$ where $x(z)=[1,z,z^{2},z^{3},z^{4}]^{\top}$. We show results for both the standard high-fidelity only and the proposed multifidelity approaches for three different computational budgets of 10, 100, and 1000.
Experiments using exact model statistics. For this analytical example, the exact model statistics $\sigma_{i}$, $\rho_{12}$, and $\Gamma_{12}$ and $\Gamma_{22}$ can be calculated. For a given budget, we use the exact statistics to compute the sample allocations according to eq. 10. The resulting sample allocation is shown in Table 1.
@@
Based on these exact statistics, the heuristic scalar control variate coefficient (9) is $\alpha=12.79$; the optimal scalar control variate coefficient is $\alpha^{*}=11.97$; and the optimal matrix control variate coefficient is
$A^{*}=\begin{pmatrix}1.8\text{e}1&-8.0\text{e}0&2.2\text{e}0&-2.7\text{e-}1&2.4\text{e-}2\\ 2.2\text{e}2&-1.1\text{e}2&2.7\text{e}1&-3.6\text{e}0&2.9\text{e-}1\\ 2.5\text{e}3&-1.3\text{e}3&3.0\text{e}2&-4.2\text{e}1&3.1\text{e}0\\ 2.6\text{e}4&-1.3\text{e}4&3.1\text{e}3&-4.2\text{e}2&2.9\text{e}1\\ 2.3\text{e}5&-1.1\text{e}5&2.7\text{e}4&-3.7\text{e}3&2.4\text{e}2\end{pmatrix}.$
To facilitate comparison between the standard high-fidelity only (HF) training approach and our proposed multifidelity (MF) training approach, we first focus on multifidelity results using the optimal scalar coefficient $\alpha^{*}$, which is our general recommendation. We repeat the model learning process over 500 independent realizations of training data to analyze robustness of the proposed method to variations in training data. In Figure 1, the left subplot plots realizations of the first element of $\hat{c}_{XY}$, the central subplot realizations of the first regression coefficient $\hat{\beta}_{1}$, and the right subplot realizations of the regression model prediction at $z=5$, $\hat{f}(z=5;\hat{\beta})$. The replicate mean over the 500 training realizations is also shown, along with the truth that would be obtained from taking exact expectations. The realizations are plotted with some transparency so that regions with more realizations appear more densely colored. These numerical results illustrate that the multi-fidelity estimators we have introduced are unbiased (Theorem 3.1), and that the multifidelity estimators exhibit lower variance than their high-fidelity counterparts at all computational budgets.
@@
Figure 1: Exponential function example: 500 realizations (semi-transparent colored lines) of estimators for first entry of $\hat{c}_{XY}$ (left), first regression coefficient (center), and regression model prediction $\hat{f}(z=5;\hat{\beta})$ (right), when true model statistics are known. Black lines denote true values and lines with markers show the mean over the 500 realizations of training data.
Figure 2 illustrates that the variance reduction achieved by our multifidelity training approach yields more accurate and robust learned models. Figure 2(a) plots realizations of the fitted regression models $\hat{f}(z;\hat{\beta})$ using both the standard high-fidelity only (HF) and multifidelity (MF) training approaches for a computational budget of 100, together with a comparison to the exact regression model $\hat{f}(z;\beta^{*})$, which can be computed for this analytical example. Note that both HF and MF learned models are unbiased estimators for the exact regression model, but that the HF approach exhibits higher variance and is less robust to variations of the training data, even with the relatively high budget of 100 high-fidelity samples for this simple example. Figure 2(b) plots the generalization error of the learned regression models trained with the HF and MF approaches at different computational budgets. For this experiment, 1000 random test data are generated from the input distribution for each learned model replicate, and the mean relative error of the learned model, $\left|\frac{\hat{f}(z;\hat{\beta})-f^{(1)}(z)}{f^{(1)}(z)}\right|$, is evaluated for 500 realizations of the HF and MF learned models. The mean and first standard deviations over these 500 realizations are shown. For this example, the MF training has a similar mean generalization error than the standard HF-only training using the same computational budget, but a significantly reduced variance: the high-fidelity only training can lead to much higher generalization errors at low computational budgets. This illustrates the power of our multifidelity training strategy to learn more robust models from scarce high-fidelity data.
(a) Visualization of 500 HF and MF learned models trained from independent realizations of training data with a computational budget of 100.
(b) Generalization error over 1000 unseen test data. Plotted lines and shaded regions are the mean and first standard deviation over 500 independently trained models.
Figure 2: Exponential example: comparing learned models learned with the standard high-fidelity only (HF) and proposed multifidelity (MF) training approach.
Experiments using inexact model statistics. We now explore the effect of estimating the statistics required to set the control variate coefficient $A$ and determine the model allocation in our multifidelity approach, and compare the performance of different control variate coefficient choices when these statistic estimates vary in quality. Figure 3 plots the (generalized) variance of estimates of $\hat{c}_{XY}$ (top row), $\hat{\beta}$ (middle row), and $\hat{f}(z;\hat{\beta})$ (bottom row) over 500 realizations of training data. In each subplot, the standard high-fidelity only approach is compared to the proposed multifidelity approach with the three choices of control variate coefficient discussed in Section 3.2.1. The columns of Figure 3 correspond to using statistics estimated exactly (left), and using 100 (middle), and 10 (right) pilot samples. Estimated statistics are used to determine both the control variate coefficient and the model allocation: we find that using estimated model statistics to determine the sample allocation according to (10) yields a 1% (5%) variation in $m_{1}$ and a 3% (25%) variation in $m_{2}$ when 100 (10) pilot samples are used.
Figure 3: Analytical example: convergence of multifidelity linear regression estimators for $\hat{c}_{XY}$ (top), $\hat{\beta}$ (middle), and $\hat{f}(z;\hat{\beta})$ (bottom) when model statistics are exact (left), estimated using 100 pilot samples (center), or 10 pilot samples (right).
Our results in Figure 3 show that for this analytical example, the multifidelity estimators generally have lower variance than their standard single-fidelity counterparts, which leads to more robust models with lower average errors, as discussed above. These results therefore demonstrate that our proposed multifidelity training approach consistently leads to more robust and accurate learned models for this analytical example. Figure 3 also illustrates our theoretical results from Section 3.3 concerning the optimal choices of control variate coefficient: when exact statistics are known, the optimal matrix choice exhibits the greatest variance reduction, while the scalar choices, MF-$\alpha^{*}$ and MF-$\alpha^{\rm mean}$, exhibit similar performance. For this example, the optimal matrix choice remains the most effective even when statistics are estimated from 100 pilot samples, but when statistics are estimated using just 10 pilot samples this ‘optimal’ choice leads to sub-optimal variance reduction in the regressor predictions and in fact higher variance in the learned parameters than the high-fidelity only training achieves. This underscores our general recommendation to use the optimal scalar control variate coefficient in practical settings where the model statistics are unknown and must be estimated, although we note that for this example the heuristic scalar choice has similar performance. In fact, the heuristic choice slightly outperforms the ‘optimal’ scalar choice when model statistics are estimated from 10 pilot samples for this analytical example, demonstrating that this may also be a reasonable choice in practice. With either scalar control variate coefficient choice, the multifidelity models trained with the lowest budget (just 8 high-fidelity samples) have similar prediction variance to standard learned models trained with hundreds of high-fidelity samples, demonstrating the efficacy of the multifidelity approach for learning robust and accurate models in the scarce data regime.
@@ 4.2.1 Convection-diffusion-reaction (CDR) problem
$\begin{split}\frac{\partial u}{\partial t}&=\kappa\Delta u-U_{\rm vel}\nabla u+s(u,z)\\ s_{i}(u,z)&=\gamma_{i}\nu_{i}\left(\frac{W_{i}}{\rho_{\text{mix}}}\right)\left(\frac{\rho_{\text{mix}}Y_{F}}{W_{F}}\right)^{\nu_{F}}\left(\frac{\rho_{\text{mix}}Y_{O}}{W_{O}}\right)^{\nu_{O}}A_{\text{pe}}\exp\left(-\frac{E}{RT}\right),\,\quad\text{for }i=F,O,P,\\ s_{T}(u,z)&=s_{P}(u,z)Q.\end{split}$ (22)
In (22), the thermo-chemical state is $u(z,t)=[Y_{F},Y_{O},Y_{P},T]^{\top}$, where $Y_{F}$, $Y_{O}$, and $Y_{P}$ are the mass fractions of the fuel, oxidizer, and product, respectively, and $T$ is the temperature; $\kappa$ is the molecular diffusivity, $U_{\rm vel}$ is the velocity field, and $s$ is the nonlinear reaction source term. Additionally, $\gamma_{P}=1$ and $\gamma_{O}=\gamma_{F}=-1$. The reaction modeled is a one-step hydrogen combustion given by $2\text{H}_{2}+\text{O}_{2}\rightarrow 2\text{H}_{2}\text{O}$, where hydrogen is the fuel, oxygen acts as the oxidizer, and water is the product. The source terms are defined using $A_{\text{pe}}$ as the pre-exponential factor of the Arrhenius equation, $E$ as the activation energy, $W_{i}$ as the molecular weight of species $i$, $\rho_{\text{mix}}$ as the density of the mixture, $R$ as the universal gas constant, and $Q$ as the heat of the reaction.
@@
where $T_{i}$ is the temperature at the inlet, $T_{0}$ is the temperature of the left wall of the domain, and $\phi$ is the fuel:oxidizer ratio of the premixed inflow. The parameters $A_{\text{pe}}$ and $E$ are assumed to be log-uniformly distributed and $T_{i}$, $T_{0}$, and $\phi$ are assumed to be uniformly distributed. We fit a quadratic model to these five inputs, so that the feature vector $x$ has length 21 (one constant, five linear, and 15 quadratic features). Due to the varying scales of the input parameters, the parameters are scaled to be order 1 before monomial features are computed.
@@ 4.2.2 Results: CDR problem
Because analytical model statistics are not easily obtainable for this example, we compute ‘exact’ reference statistics using all $2.4\times 10^{5}$ samples in the data set from which we bootstrap the training data. Figure 4 (analogously to Figure 3) plots the (generalized) variance of estimates of $\hat{c}_{XY}$ (top row), $\hat{\beta}$ (middle row), and $\hat{f}(z;\hat{\beta})$ (bottom row) over 500 realizations of training data for computational budgets of 10, 100, and 1000. The columns of Figure 4 correspond to using all $2.4\times 10^{5}$ samples (left), 100 pilot samples (center), and 10 pilot samples (right) to estimate model statistics that are then used to determine the control variate coefficients and sample allocations in our multifidelity approaches. Table 3 provides the sample allocation obtained using all $2.4\times 10^{5}$ samples. We find that using fewer pilot samples to determine the sample allocation according to (10) yields a 1% (10%) variation in $m_{1}$ and a 6% (30%) variation in $m_{2}$ when 100 (10) pilot samples are used.
@@
Table 3: Sample allocation based on (10) for the CDR problem based on reference model statistics computed using all available samples in the data set.
Figure 4: PDE model problem: convergence of multifidelity linear regression estimators for $\hat{c}_{XY}$ (top), $\hat{\beta}$ (middle), and $\hat{f}(z;\hat{\beta})$ (bottom), when model statistics are estimated using $10^{5}$ (left), 100 (center), or 10 (right) pilot samples. Results for the multifidelity approach with the optimal matrix coefficient are omitted in the second and third columns because their variances are so large that they would significantly distort the plot axes.
For this PDE model problem, Figure 4 shows that when the model statistics are estimated using many samples, the optimal matrix control variate coefficient leads to the greatest variance reduction. However, when only 100 or 10 pilot samples are available, the ‘optimal’ coefficient actually leads to much higher multifidelity variance than the standard high-fidelity only approach, so we have chosen to omit the MF-$A^{*}$ lines from the second and third columns because they would distort the axes significantly. However, both choices of scalar control variate coefficient consistently lead to multifidelity linear regression estimators with lower variance than their high-fidelity counterparts, and the optimal scalar coefficient consistently performs better than the heuristic choice. For this reason, our general recommendation is to use the optimal scalar choice, although for complex engineering problems it may be easier to use the heuristic choice, and our numerical results illustrate that this still leads to practical gains: for this problem, both scalar coefficient choices yield multifidelity learned models that, when trained with the lowest computational budget (just 4 high-fidelity samples), achieve a similar average error to standard learned models trained with $\mathcal{O}(10-100)$ times as many high-fidelity samples.
Finally, we plot in Figure 5 the mean generalization error of the HF and MF learned models over 1000 test data bootstrapped independently from the available data set. The mean and first standard deviation over 500 learned models trained on independent realizations of training and test data are shown. The multifidelity results use the optimal scalar coefficient, our general recommendation. Figure 5 shows that our MF training approach achieves a mean relative error that is several times less than that of the standard HF only training approach. The error reduction is especially stark at our lowest computational budget, again demonstrating the efficacy of the multifidelity approach for learning robust and accurate models in the scarce data regime.
Figure 5: Generalization error over 1000 unseen test data. Plotted lines and shaded regions are the mean and first standard deviation over 500 learned models trained on independently realizations of training data.